# DPO: Dual-Perturbation Optimization for Test-time Adaptation in 3D Object Detection

### v1

@@ Abstract.
LiDAR-based 3D object detection has seen impressive advances in recent times. However, deploying trained 3D detectors in the real world often yields unsatisfactory performance when the distribution of the test data significantly deviates from the training data due to different weather conditions, object sizes, etc. A key factor in this performance degradation is the diminished generalizability of pre-trained models, which creates a sharp loss landscape during training. Such sharpness, when encountered during testing, can precipitate significant performance declines, even with minor data variations. To address the aforementioned challenges, we propose dual-perturbation optimization (DPO) for Test-time Adaptation in 3D Object Detection (TTA-3OD). We minimize the sharpness to cultivate a flat loss landscape to ensure model resiliency to minor data variations, thereby enhancing the generalization of the adaptation process. To fully capture the inherent variability of the test point clouds, we further introduce adversarial perturbation to the input BEV features to better simulate the noisy test environment. As the dual perturbation strategy relies on trustworthy supervision signals, we utilize a reliable Hungarian matcher to filter out pseudo-labels sensitive to perturbations. Additionally, we introduce early Hungarian cutoff to avoid error accumulation from incorrect pseudo-labels by halting the adaptation process. Extensive experiments across three types of transfer tasks demonstrate that the proposed DPO significantly surpasses previous state-of-the-art approaches, specifically on Waymo $\rightarrow$ KITTI, outperforming the most competitive baseline by 57.72% in $\text{AP}_{\text{3D}}$ and reaching 91% of the fully supervised upper bound. Our code is available at: https://github.com/Jo-wang/DPO.
@@ 1. Introduction
LiDAR-based 3D object detection has gained significant attention with the rapid advancements in autonomous driving (Li et al., 2021; Deng et al., 2021; Qian et al., 2022; Wang et al., 2019; Arnold et al., 2019; Li et al., 2019; Ju et al., 2022) and robotics (Montes et al., 2020; Zhou et al., 2022), where mainstream 3D detectors are developed to interpret pure point clouds or fuse multimodal knowledge, commonly incorporating camera images (Wang et al., 2023a; Liu et al., 2023). However, deploying either point clouds-based or multimodal 3D detection models in real-world scenarios often leads to performance degradation due to distribution shifts between the training data and the encountered real-world data. For instance, a 3D detector trained on the nuScenes dataset (Caesar et al., 2020) might suffer a performance drop when applied to the KITTI dataset (Geiger et al., 2012) due to variations in object sizes and the number of beams. This is known as cross-dataset shift. Additionally, the shift can arise from real-world disturbances, termed as corruption-based shift (Hahner et al., 2022; Kong et al., 2023; Dong et al., 2023), which includes challenges like diverse weather conditions and sensor malfunctions. Moreover, multiple factors are likely to be concurrent, for instance, deploying a 3D detector in a different city while suffering severe snow. This scenario is termed as composite domain shift.
@@
Domain adaptive 3D object detection has been discovered to mitigate the performance gap brought by various domain shifts. This involves aligning features between the labeled training data and the shifted test data to learn a domain-invariant representation (Luo et al., 2021; Zhang et al., 2021; Zeng et al., 2021) or conducting self-training with the aid of selected pseudo-labels (Chen et al., 2023; Yang et al., 2021, 2022; Peng et al., 2023; Li et al., 2023a). However, these approaches necessitate extensive training over multiple epochs on both training and test sets, rendering them impractical for adaptation to the streaming data. Moreover, the exposure of the training data can significantly compromise its privacy, especially when it contains sensitive user information (e.g., user vehicle trajectories and individuals).
@@
To bridge the performance gap induced by domain shifts, while safeguarding the training data privacy and enabling swift adaptation, test-time adaptation (TTA) emerges as an ideal solution. Prior research on TTA typically adapts a source pre-trained model to the unlabeled test data, either through updating a selected subset of parameters (e.g., BatchNorm layers) (Wang et al., 2021; Niu et al., 2023; Sun et al., 2023), or employing the mean-teacher model (Wang et al., 2022, 2023c; Yuan et al., 2023; Tomar et al., 2023a) within a single epoch. However, these TTA works currently applied in image classification are inadequate for addressing the dual demands (i.e., object localization and classification) for supervision signals inherent in detection tasks. Within this context, MemCLR (VS et al., 2023) stands out by refining the Region of Interest (RoI) features of detected objects through a transformer-based memory module for 2D object detection. Nevertheless, the stored target representations derived from the source pre-trained model cause performance degradation due to distribution shifts. These limitations pose significant challenges in utilizing previous TTA techniques for 3D object detection.
@@ 2.1. Domain Adaptive 3D Object Detection
Adaptation for 3D Object Detection focuses on transferring knowledge from 3D detectors trained on labeled source point clouds to unlabeled target domains, effectively reducing the domain discrepancies across diverse 3D environments such as variations in object statistics (Wang et al., 2020; Tsai et al., 2023b), weather conditions (Xu et al., 2021; Hahner et al., 2022), sensor differences (Rist et al., 2019; Gu et al., 2021; Wei et al., 2022), sensor failures (Kong et al., 2023), and the synthetic-to-real gap (Saleh et al., 2019; DeBortoli et al., 2021; Lehner et al., 2022). Strategies to overcome these challenges include adversarial feature alignment (Zhang et al., 2021), 3D pseudo-labels (Yang et al., 2021, 2022; Chen et al., 2023; Saltori et al., 2020; You et al., 2022; Li et al., 2023b; Peng et al., 2023; Huang et al., 2024; Tsai et al., 2023a), the mean-teacher model (Luo et al., 2021; Hegde et al., 2023) for prediction consistency, and contrastive learning (Zeng et al., 2021). Nonetheless, these cross-domain adaptation methods typically necessitate adaptation over multiple epochs, making them less suited for real-time test scenarios.
@@ 2.2. Test-time Adaptation in 2D Vision Tasks
Test-time adaptation (TTA) (Liang et al., 2023; Wang et al., 2023b) is designed to address domain shifts between the training and testing data (Wang et al., 2023b) during inference time. As a representative, Tent (Wang et al., 2021) leverages entropy minimization for BatchNorm adaptation. Subsequent works (Hong et al., 2023; Gong et al., 2022; Song et al., 2023; Nguyen et al., 2023; Niloy et al., 2024) such as EATA (Niu et al., 2022), identifies reliable and nonredundant samples to optimize. DUA (Mirza et al., 2022) introduces adaptive momentum in a new normalization layer whereas RoTTA (Yuan et al., 2023) and DELTA (Zhao et al., 2023) leverage global statistics for batch norm updates. Furthermore, SoTTA (Gong et al., 2024) and SAR (Niu et al., 2023) improve BatchNorm optimization by minimizing the loss sharpness. Alternatively, some approaches optimize the entire network through the mean-teacher framework for stable supervision (Wang et al., 2022; Tomar et al., 2023b), generate reliable pseudo-labels for self-training (Goyal et al., 2022; Zeng et al., 2024), employ feature clustering (Chen et al., 2022; Jung et al., 2023; Wang et al., 2023c), and utilizing augmentations to enhance model robustness (Zhang et al., 2022). However, these TTA methods are developed for general image classification. Additionally, MemCLR (VS et al., 2023) applies TTA for image-based 2D object detection, using a mean-teacher approach to align instance-level features. Nevertheless, the applicability of these image-based TTA methods to object detection from 3D point clouds remains unexplored.
@@ 3.2. Minimizing Sharpness in the Weight Space
The sharpness of the training loss, is the rate of change in the surrounding region of the loss landscape. It has been identified to be empirically correlated with the generalization error (Ghadimi and Lan, 2013; Jiang et al., 2020; Keskar et al., 2017). Motivated by this, recent works propose to reduce the loss sharpness during the training phase, aiming to improve the generalization capabilities of the model. One notable example is Sharpness-Aware Minimization (SAM), which enhances model training by integrating and optimizing the worst-case perturbations in model weights. The fundamental principle of SAM is that by minimizing the loss with respect to maximally perturbed weights within a vicinity, the entire vicinity (i.e., all losses within it) is minimized. This directs the optimization trajectory toward a flat minima in the loss landscape. A flat minima is indicative of superior generalization capabilities, as the loss over it is less sensitive to large perturbations and/or noise in the model weights, unlike sharp minima. In the context of TTA-3OD, the loss sharpness (Andriushchenko and Flammarion, 2022) during the adaptation can be formally defined as follows:
@@
To expedite computation, we omit the second-order terms.
Figure 1. (1) Loss contour for weight perturbation $\hat{\epsilon}_{w}$ (left); (2) loss profile view for input perturbation $\hat{\epsilon}_{z}$ (right). Our goal is to optimize the loss towards flat minima while ensuring the model’s resilience to data perturbations. Darker colors indicate lower loss values.
While this improves model generalization by targeting non-sharp minima within the loss landscape, the optimized perturbations to weights do not directly deal with variations and/or noise of the input test scenes. When facing the test data, the detection performance is substantially degraded due to the severe data-level corruptions in the test point cloud. For example, when the 3D detection model suffers heavy snow and shifted object scales, simultaneously, the performance drops from 73.45% to 3.84% in $\text{AP}_{\text{3D}}$. The empirical evidence suggests that augmenting the baseline model with SAM (Niu et al., 2023) results in a marginal improvement of only 0.9% in $\text{AP}_{\text{3D}}$, indicating its ineffectiveness in bridging the domain gap in the test 3D scenes.
@@ 4.1. Experimental Setup
We leverage the OpenPCDet framework (Team, 2020). Experiments are conducted on a single NVIDIA RTX A6000 GPU with 48 GB of memory. We opt for a batch size of $8$ and fix the hyperparameters $\rho=1e-4,\alpha=0.08,\gamma=0.5,\eta=10^{-3}$. For evaluation purposes, we adhere to the official metrics of the KITTI benchmark, reporting the average precision for the car class [^1] The experiments of other classes are shown in the supplementary material. in both 3D (i.e., $\text{AP}_{\text{3D}}$) and bird’s eye view (i.e., $\text{AP}_{\text{BEV}}$) perspectives, calculated over 40 recall positions and applying a $0.7$ IoU threshold. The closed gap (Yang et al., 2021) is calculated as: $\frac{\mathrm{AP}_{\text{method }}-\mathrm{AP}_{\text{No Adapt. }}}{\mathrm{AP}_{\text{Oracle }}-\mathrm{AP}_{\text{No Adapt. }}}\times 100\%$.
@@
  • SN (Wang et al., 2020): a weakly supervised domain adaptation method for 3D detection that adjusts source object sizes using target domain statistics for pre-training.
@@
  • CoTTA (Wang et al., 2022): a TTA strategy that leverages mean-teacher framework to provide supervisory signals through augmentations and employs random neuron restoration to retain source knowledge.
@@ 4.2. Experimental Results
To assess DPO’s efficacy against corruption-induced shifts, we performed experiments on KITTI $\rightarrow$ KITTI-C with heavy severity of eight real-world corruption, evaluating $\text{AP}_{\text{3D}}$ in hard difficulty scenarios. As indicated in Table 2, DPO outperforms all TTA baselines in terms of Mean $\text{AP}_{\text{3D}}$, exceeding the performance of the closest competitive baseline, Tent, by 1.2%. DPO consistently achieves top performance across most corruption types, demonstrating the enhanced robustness of DPO and its effectiveness in adapting 3D models to a wide array of corrupted environments.
@@ 4.3. Parameter Sensitivity
4.3.2. Pseudo-label Threshold $\alpha$.
The pseudo-label threshold $\alpha$ exhibits a similar pattern for $\text{AP}_{\text{BEV}}$, remaining relatively stable across different thresholds. However, setting the threshold too low (i.e., 0.04) results in a performance drop in terms of $\text{AP}_{\text{3D}}$ due to too few pseudo-labeled 3D boxes being selected to update model weights. This highlights the importance of choosing an appropriate proportion of pseudo-labels for adaptation. For $\alpha$ values within the 0.08 to 0.20 range, the performance in $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ exhibit stability, evidenced by maximum fluctuations of 0.83 and 0.07, respectively. This stability underscores the robustness of the selected threshold.
@@ 4.4. Ablation Study
To gain an in-depth understanding of how individual components of DPO contribute to overall performance improvements, we conduct an ablation study by incrementally integrating each component into adaptation and assess the resulting performance on the nuScenes $\rightarrow$ KITTI task. Table 3 presents the impact of the proposed components across three difficulty levels of the KITTI dataset in terms of AP score. Here, Pert. $\Theta_{t}$ represents weight space perturbation, Pert. $Z_{t}$ denotes input perturbation, and Matcher refers to the Hungarian Matching mechanism for pseudo-label selection. Compared to the self-training baseline (row 1), introducing a single weight space perturbation (row 2) markedly boosts $\text{AP}_{\text{BEV}}$, but shows a negative effect in $\text{AP}_{\text{3D}}$ (58.78 $\rightarrow$ 56.36 at easy difficulty), which indicates the limitations of direct using SAM for the TTA-3OD task. While, further incorporating the proposed input perturbation (row 3) and optimizing the detection model with a dual-perturbation, gains significant performance, elevating $\text{AP}_{BEV}$ and $\text{AP}_{\text{3D}}$ for all levels of difficulty. The Hungarian matcher, ensuring reliable pseudo-label selection even with single weight perturbation, is evidenced by non-neglectable gains over solely applying weight perturbation (rows 2, 4). Comprehensive utilization of our proposed DPO secures the highest performance for both $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ across all difficulty tiers.
@@
Figure 5. Performance trend and variation in the number of point clouds for weight updates across different thresholds $C_{\text{stop}}$.
To assess the necessity of employing the early Hungarian cutoff, we examined its sensitivity on the nuScenes $\rightarrow$ Waymo task, with results depicted in Figure 5. When the moving-average Hungarian cost, as determined by Eqn. (12), falls below a specified threshold, for instance, $C_{\text{stop}}=4.8$, the model updates its parameter weights through self-training on the first 48 test point clouds. Subsequently, it infers the remaining point clouds directly, bypassing further self-training. There is an obvious trend that, during the self-training, the moving-average Hungarian cost converges very fast. For example, training the first 64 test samples induces a decrease in cost from 4.8 to 3.8, and results in a substantial performance enhancement (9.37 in $\text{AP}_{\text{3D}}$). Conversely, a reduction in cost from 3.6 to 3.4, involving a significantly larger computational effort (training on 144 test samples), yields only a marginal improvement in performance (3.53 in $\text{AP}_{\text{3D}}$). The reason is that the accumulation of errors in pseudo-labels hinders further increasing the adaption performance. These observations highlight the importance of the proposed Hungarian cost-based early stopping mechanism, which utilizes only a small proportion of test batches to maximize the adaption performance without computational burden.
4.4.3. Impact of Updating Strategy
@@
Table 4. BatchNorm v.s. full model updating strategies.
@@
Table 5. Results of Waymo $\rightarrow$ KITTI using PV-RCNN (Shi et al., 2020).

### v2

@@ Abstract.
LiDAR-based 3D object detection has seen impressive advances in recent times. However, deploying trained 3D detectors in the real world often yields unsatisfactory performance when the distribution of the test data significantly deviates from the training data due to different weather conditions, object sizes, etc. A key factor in this performance degradation is the diminished generalizability of pre-trained models, which creates a sharp loss landscape during training. Such sharpness, when encountered during testing, can precipitate significant performance declines, even with minor data variations. To address the aforementioned challenges, we propose dual-perturbation optimization (DPO) for Test-time Adaptation in 3D Object Detection (TTA-3OD). We minimize the sharpness to cultivate a flat loss landscape to ensure model resiliency to minor data variations, thereby enhancing the generalization of the adaptation process. To fully capture the inherent variability of the test point clouds, we further introduce adversarial perturbation to the input BEV features to better simulate the noisy test environment. As the dual perturbation strategy relies on trustworthy supervision signals, we utilize a reliable Hungarian matcher to filter out pseudo-labels sensitive to perturbations. Additionally, we introduce early Hungarian cutoff to avoid error accumulation from incorrect pseudo-labels by halting the adaptation process. Extensive experiments across three types of transfer tasks demonstrate that the proposed DPO significantly surpasses previous state-of-the-art approaches, specifically on Waymo $\rightarrow$ KITTI, outperforming the most competitive baseline by 57.72% in $\text{AP}_{\text{3D}}$ and reaching 91% of the fully supervised upper bound. Our code is available at https://github.com/Jo-wang/DPO.
Test-time Adaptation, 3D Object Detection.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/XXXXXX.XXXXXXjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681040isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Scene understandingccs: Computing methodologies Online learning settingsccs: Computing methodologies Object detectionccs: Computing methodologies Transfer learning
@@ 1. Introduction
LiDAR-based 3D object detection has gained significant attention with the rapid advancements in autonomous driving (Li et al., 2021; Deng et al., 2021; Luo et al., 2023a; Qian et al., 2022; Luo et al., 2023b; Ju et al., 2022) and robotics (Montes et al., 2020; Zhou et al., 2022), where mainstream 3D detectors are developed to interpret pure point clouds or fuse multimodal knowledge, commonly incorporating camera images (Wang et al., 2023a; Chen et al., 2021b; Liu et al., 2023). However, deploying either point clouds-based or multimodal 3D detection models in real-world scenarios often leads to performance degradation due to distribution shifts between the training data and the encountered real-world data. For instance, a 3D detector trained on the nuScenes dataset (Caesar et al., 2020) might suffer a performance drop when applied to the KITTI dataset (Geiger et al., 2012) due to variations in object sizes and the number of beams. This is known as cross-dataset shift. Additionally, the shift can arise from real-world disturbances, termed as corruption-based shift (Hahner et al., 2022; Kong et al., 2023; Dong et al., 2023), which includes challenges like diverse weather conditions and sensor malfunctions. Moreover, multiple factors are likely to be concurrent, for instance, deploying a 3D detector in a different city while suffering severe snow. This scenario is termed as composite domain shift.
@@
Domain adaption has been discovered (Wang et al., 2022b, 2023b; Luo et al., 2023c; Chen et al., 2021a) to mitigate the performance gap brought by various domain shifts. In 3D object detection, this involves aligning features between the labeled training data and the shifted test data to learn a domain-invariant representation (Luo et al., 2021; Zhang et al., 2021; Zeng et al., 2021) or conducting self-training with the aid of selected pseudo-labels (Chen et al., 2023; Yang et al., 2021, 2022; Peng et al., 2023; Li et al., 2023a). However, these approaches necessitate extensive training over multiple epochs on both training and test sets, rendering them impractical for adaptation to the streaming data. Moreover, the exposure of the training data can significantly compromise its privacy, especially when it contains sensitive user information (e.g., user vehicle trajectories and individuals).
@@
To bridge the performance gap induced by domain shifts, while safeguarding the training data privacy and enabling swift adaptation, test-time adaptation (TTA) emerges as an ideal solution. Prior research on TTA typically adapts a source pre-trained model to the unlabeled test data, either through updating a selected subset of parameters (e.g., BatchNorm layers) (Wang et al., 2021; Niu et al., 2023; Sun et al., 2023), or employing the mean-teacher model (Wang et al., 2022a, 2023d; Yuan et al., 2023; Tomar et al., 2023a) within a single epoch. However, these TTA works currently applied in image classification are inadequate for addressing the dual demands (i.e., object localization and classification) for supervision signals inherent in detection tasks. Within this context, MemCLR (VS et al., 2023) stands out by refining the Region of Interest (RoI) features of detected objects through a transformer-based memory module for 2D object detection. Nevertheless, the stored target representations derived from the source pre-trained model cause performance degradation due to distribution shifts. These limitations pose significant challenges in utilizing previous TTA techniques for 3D object detection.
@@ 2.1. Domain Adaptive 3D Object Detection
Adaptation for 3D Object Detection focuses on transferring knowledge from 3D detectors trained on labeled source point clouds to unlabeled target domains, effectively reducing the domain discrepancies across diverse 3D environments such as variations in object statistics (Wang et al., 2020; Tsai et al., 2023b), weather conditions (Xu et al., 2021; Hahner et al., 2022), sensor differences (Rist et al., 2019; Gu et al., 2021; Wei et al., 2022), sensor failures (Kong et al., 2023), and the synthetic-to-real gap (Saleh et al., 2019; DeBortoli et al., 2021; Lehner et al., 2022). Strategies to overcome these challenges include adversarial feature alignment (Zhang et al., 2021), 3D pseudo-labels (Yang et al., 2021, 2022; Chen et al., 2023; Saltori et al., 2020; Chen et al., 2024; You et al., 2022; Li et al., 2023b; Peng et al., 2023; Huang et al., 2024; Tsai et al., 2023a), the mean-teacher model (Luo et al., 2021; Hegde et al., 2023) for prediction consistency, and contrastive learning (Zeng et al., 2021). Nonetheless, these cross-domain adaptation methods typically necessitate adaptation over multiple epochs, making them less suited for real-time test scenarios.
@@ 2.2. Test-time Adaptation in 2D Vision Tasks
Test-time adaptation (TTA) (Liang et al., 2023; Wang et al., 2023c) is designed to address domain shifts between the training and testing data (Wang et al., 2023c) during inference time. As a representative, Tent (Wang et al., 2021) leverages entropy minimization for BatchNorm adaptation. Subsequent works (Hong et al., 2023; Gong et al., 2022; Song et al., 2023; Nguyen et al., 2023; Niloy et al., 2024) such as EATA (Niu et al., 2022), identifies reliable and nonredundant samples to optimize. DUA (Mirza et al., 2022) introduces adaptive momentum in a new normalization layer whereas RoTTA (Yuan et al., 2023) and DELTA (Zhao et al., 2023) leverage global statistics for batch norm updates. Furthermore, SoTTA (Gong et al., 2024) and SAR (Niu et al., 2023) improve BatchNorm optimization by minimizing the loss sharpness. Alternatively, some approaches optimize the entire network through the mean-teacher framework for stable supervision (Wang et al., 2022a; Tomar et al., 2023b), generate reliable pseudo-labels for self-training (Goyal et al., 2022; Zeng et al., 2024), employ feature clustering (Chen et al., 2022; Jung et al., 2023; Wang et al., 2023d), and utilizing augmentations to enhance model robustness (Zhang et al., 2022). However, these TTA methods are developed for general image classification. Additionally, MemCLR (VS et al., 2023) applies TTA for image-based 2D object detection, using a mean-teacher approach to align instance-level features. Nevertheless, the applicability of these image-based TTA methods to object detection from 3D point clouds remains unexplored.
@@ 3.2. Minimizing Sharpness in the Weight Space
The sharpness of the training loss, is the rate of change in the surrounding region of the loss landscape. It has been identified to be empirically correlated with the generalization error (Ghadimi and Lan, 2013; Jiang et al., 2020; Keskar et al., 2017). Motivated by this, recent works propose to reduce the loss sharpness during the training phase, aiming to improve the generalization capabilities of the model. One notable example is Sharpness-Aware Minimization (SAM), which enhances model training by integrating and optimizing the worst-case perturbations in model weights. The fundamental principle of SAM is that by minimizing the loss with respect to maximally perturbed weights within a vicinity, the entire vicinity (i.e., all losses within it) is minimized. This directs the optimization trajectory toward a flat minima in the loss landscape. A flat minima is indicative of superior generalization capabilities, as the loss over it is less sensitive to large perturbations and/or noise in the model weights, unlike sharp minima.
In the context of TTA-3OD, the loss sharpness (Andriushchenko and Flammarion, 2022) during the adaptation can be formally defined as follows:
@@
To expedite computation, we omit the second-order term.
Figure 1. (1) Loss contour for weight perturbation $\hat{\epsilon}_{w}$ (left); (2) The loss profile view for input perturbation $\hat{\epsilon}_{z}$ (right). Our goal is to optimize the loss towards flat minima while ensuring the model’s resilience to data perturbations. Darker colors indicate lower loss values.
While this improves model generalization by targeting non-sharp minima within the loss landscape, the optimized perturbations to weights do not directly deal with variations and/or noise of the input test scenes. When facing the test data, the detection performance is substantially degraded due to the severe data-level corruptions in the test point cloud. For example, when suffering heavy snow and shifted object scales, simultaneously, the performance drops from 73.45% to 3.84% in $\text{AP}_{\text{3D}}$. The empirical evidence suggests that augmenting the baseline model with SAM (Niu et al., 2023) results in a marginal improvement of only 0.9% in $\text{AP}_{\text{3D}}$, indicating its ineffectiveness in bridging the domain gap in the test 3D scenes.
@@ 4.1. Experimental Setup
We leverage the OpenPCDet framework (Team, 2020). Experiments are conducted on a single NVIDIA RTX A6000 GPU with 48 GB of memory. We opt for a batch size of $8$ and fix the hyperparameters $\rho=1e-4,\alpha=0.08,\gamma=0.5,\eta=10^{-3}$. For evaluation purposes, we adhere to the official metrics of the KITTI benchmark, reporting the average precision for the car class in both 3D (i.e., $\text{AP}_{\text{3D}}$) and bird’s eye view (i.e., $\text{AP}_{\text{BEV}}$) perspectives, calculated over 40 recall positions and applying a $0.7$ IoU threshold. The closed gap (Yang et al., 2021) is calculated as: $\frac{\mathrm{AP}_{\text{method }}-\mathrm{AP}_{\text{No Adapt. }}}{\mathrm{AP}_{\text{Oracle }}-\mathrm{AP}_{\text{No Adapt. }}}\times 100\%$.
@@
  • SN (Wang et al., 2020): weakly supervised domain adaptive 3D detection that adjusts source object sizes using target domain statistics.
@@
  • CoTTA (Wang et al., 2022a): a TTA strategy that leverages mean-teacher framework to provide supervisory signals through augmentations and employs random neuron restoration to retain source knowledge.
@@ 4.2. Experimental Results
We evaluated DPO’s efficacy against corruption-induced shifts on KITTI $\rightarrow$ KITTI-C with heavy severity of eight real-world corruption by $\text{AP}_{\text{3D}}$ in hard difficulty scenarios. As indicated in Table 2, DPO outperforms all TTA baselines in terms of Mean $\text{AP}_{\text{3D}}$, exceeding the performance of the closest competitive baseline, Tent, by 1.2%. DPO consistently achieves top performance across most corruption types, demonstrating the enhanced robustness of DPO and its effectiveness in adapting 3D models to a wide array of corrupted environments.
@@ 4.3. Parameter Sensitivity
4.3.2. Pseudo-label Threshold $\alpha$
The pseudo-label threshold $\alpha$ shows a consistent pattern for $\text{AP}_{\text{BEV}}$, remaining stable across different values. However, a low threshold (i.e., 0.04) causes a drop in $\text{AP}_{\text{3D}}$ as too few pseudo-labeled 3D boxes are selected to update model weights. This emphasizes the need for an appropriate proportion of pseudo-labels for adaptation. For $\alpha$ values between 0.08 and 0.20, $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ remain stable, with maximum fluctuations of 0.83 and 0.07, respectively. This stability highlights the robustness of the selected threshold.
@@ 4.4. Ablation Study
To understand how individual components of DPO affect overall performance, we conduct an ablation study by incrementally adding each component to adaptation and evaluating performance on the nuScenes $\rightarrow$ KITTI task. Table 3 shows the impact of these components on the KITTI dataset at three difficulty levels, measured by AP score. Here, Pert. $\Theta_{t}$ represents weight space perturbation, Pert. $Z_{t}$ denotes input perturbation, and Matcher refers to the Hungarian Matching mechanism for pseudo-label selection. Compared to the self-training baseline (row 1), adding weight space perturbation (row 2) significantly improves $\text{AP}_{\text{BEV}}$ but reduces $\text{AP}_{\text{3D}}$ (58.78 $\rightarrow$ 56.36 at easy difficulty), indicating limitations of SAM for the TTA-3OD task. Incorporating input perturbation (row 3) and using dual perturbation improves performance, increasing $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ across all difficulty levels. The Hungarian matcher enhances pseudo-label selection with weight perturbation alone, as shown by the performance gains over weight perturbation alone (rows 2, 4). Using all proposed DPO components yields the highest performance for both $\text{AP}_{\text{BEV}}$ and $\text{AP}_{\text{3D}}$ across all difficulty levels.
@@
Figure 5. Performance trend and variation in the number of point clouds for weight updates across different early-stopping thresholds $C_{\text{stop}}$.
We examine the effectiveness of early Hungarian cutoff on the nuScenes $\rightarrow$ Waymo task in Figure 5. When the Hungarian cost in Eqn. (12) falls below a specified threshold, e.g., $C_{\text{stop}}=4.8$, the model updates its weights using self-training on the first 48 test point clouds and then infers the remaining point clouds directly, skipping further self-training. The moving-average Hungarian cost converges rapidly during self-training. For instance, using the first 64 test samples reduces the cost from 4.8 to 3.8 and significantly improves performance (9.37 in $\text{AP}_{\text{3D}}$). In contrast, reducing the cost from 3.6 to 3.4 with 144 test samples only marginally improves performance (3.53 in $\text{AP}_{\text{3D}}$) due to error accumulation in pseudo-labels. These results underscore the value of the Hungarian cost-based early stopping mechanism, which leverages a small portion of test batches to enhance performance without excessive computational cost.
4.4.3. Impact of Updating Strategies.
@@
Table 5. DPO results of Waymo $\rightarrow$ KITTI using PV-RCNN.